# Modelling the evolution and influence of dust in cosmological simulations that include the cold phase of the interstellar medium

### v1

@@ Abstract
While marginal in mass terms, dust grains play an outsized role in both the physics and observation of the interstellar medium (ISM). However, explicit modelling of this ISM constituent remains uncommon in large cosmological simulations. In this work, we present a model for the life-cycle of dust in the ISM that couples to the forthcoming COLIBRE galaxy formation model, which explicitly simulates the cold ISM. We follow 6 distinct grain types: 3 chemical species, including carbon and two silicate grains, with 2 size bins each. Our dust model accounts for seeding of grains from stellar ejecta, self-consistent element-by-element metal yields and depletion, grain size transfer and destruction of dust in the ISM. We detail the calibration of this model, particularly the use of a clumping factor, to account for unresolved gas clouds in which dust readily evolves. We present a fiducial run in a 253 cMpc3 cosmological volume that displays good agreement with observations of the cosmic evolution of dust density, as well as the $z=0$ galaxy dust mass function and dust scaling relations. We highlight known tensions between observational datasets of the dust-to-gas ratio as a function of metallicity depending on which metallicity calibrator is used; our model favours higher-normalisation metallicity calibrators, which agree with the observations within 0.1 dex for stellar masses $>10^{9}\;{\rm M_{\odot}}$. We compare the grain size distribution to observations of local galaxies, and find that our simulation suggests a higher concentration of small grains, associated with more diffuse ISM and the warm-neutral medium (WNM), which both play a key role in boosting H2 content. Putting these results and modelling approaches in context, we set the stage for upcoming insights into the dusty ISM of galaxies using the COLIBRE simulations.
@@ 1 Introduction
While the interaction of stars and multi-phase gas have been a primary focus in the field of galactic astrophysics, cosmic dust, a vital piece of the puzzle of how galaxies form, has received less attention. Despite comprising a marginal fraction of the baryonic mass within galaxies, these small, solid-state particulates in the interstellar medium (ISM) play a prominent role in our understanding of the physics of galaxies and their evolution. The influence of dust belies its marginal mass contribution and is felt in a variety of ISM physics; influencing cooling and heating of gas, fostering conditions for star formation, and various dynamical effects. The majority of dust is thought to comprise carbonaceous (C) grains and silicate-based grains (dominated by O, Si, Mg, Fe). Dust depletes metals from the gas phase (e.g. Jenkins, 2009; De Cia et al., 2016), catalyses molecular hydrogen (${\rm H}_{2}$) formation (Cazaux & Tielens, 2002), functions as a ‘critical coolant’, and mediates radiative transfer through absorption and re-emission of stellar radiation (e.g. Whitaker et al., 2017; Barišić et al., 2020; Galliano et al., 2021).
@@
The dust grain size distribution, composition, and spatial configuration in galaxies is critical for understanding how radiation propagates through galaxies, but remain poorly constrained due to the complex interplay between dust production/destruction mechanisms and the ISM environment. Dust forms in stellar ejecta (e.g., supernovae, AGB stars), grows via accretion from the gas phase, and is destroyed by sputtering in hot plasmas or shattering/coagulation in turbulent regions (e.g. Hirashita, 2012; Asano et al., 2013a). These processes drive a transition in dominant dust growth mechanisms with metallicity: at low metallicity, stellar ejecta dominate, while accretion becomes critical at higher metallicities ($Z\gtrsim 0.1\;Z_{\odot}$), explaining observed breaks in dust-to-gas vs. metallicity relations (e.g. Rémy-Ruyer et al., 2014; Hou et al., 2019).
Despite its importance, modelling dust evolution poses significant challenges due to additional computational complexity and profound uncertainties in physical processes like grain growth and destruction, often happening deep within dense, obscured media. Early modelling efforts focused on one-zone chemical evolution models (e.g. Dwek, 1998), but recent advances have allowed incorporation of dust into (cosmological) hydrodynamical simulations, including dust models using either a minimal grain size distribution (via e.g. a simple two-size approximation, Hirashita 2015, e.g.; Hou et al. 2019, e.g.; Granato et al. 2021, e.g.), or more extensive size distribution, modelling many grain sizes (e.g. McKinnon et al., 2018). The two-size approximation, explored in smoothed-particle hydrodynamics (SPH) and adaptive mesh refinement (AMR) codes (across idealised galaxy, zoom-in and even cosmological, hydrodynamical simulations), tracks separate populations of small ($r_{\rm grain}\lesssim 0.03\;{\rm\mu m}$) and large grains ($r_{\rm grain}\gtrsim 0.1\;{\rm\mu m}$) to capture size-dependent effects critical for modelling extinction curves and cooling rates, while forgoing excessive computational cost and the memory footprint associated with tracking a full size distribution per resolution element (Hirashita, 2015; Aoyama et al., 2017; Hou et al., 2019). However, spatial inhomogeneity of dust properties, evident in Milky Way extinction curve variations and observations of nearby galaxies (e.g. Viaene et al., 2017), and the need for multi-species treatments (carbon vs. silicate grains, e.g. Shivaei et al. 2020) remain important challenges (Granato et al., 2021; Dubois et al., 2024).
@@
Cosmological zoom-in simulations provide a means to self-consistently capture dust-gas co-evolution over cosmic timescales. These can follow how dust properties emerge with the enrichment of the ISM, avoiding the artificial initial conditions of isolated disc models (Gjergo et al., 2018; Dubois et al., 2024; Choban et al., 2024; Choban et al., 2025). While more computationally and memory intensive, zoom-in simulations lack the statistical population sampling of larger volume simulations. Large cosmological volume simulations including dust allow galaxy dust scaling relations and distributions to be tested, providing an exacting test of dust models (e.g. Dayal et al., 2010; Bekki, 2013; McKinnon et al., 2018; Li et al., 2019). Previously, volume simulations lacked the physics to explicitly trace the cold ISM, imposing an equation of state on the ISM gas, so cannot explicitly model the cool and cold ($\ll 10^{4}~{}{\rm K}$) ISM that is critically important for dust evolution processes, pushing more of the relevant physics to sub-grid modelling. Given these limitations for population scales for simulations, semi-analytic models have provided insight into the potential balance of dust evolutionary processes, and have provided a means to explore the poorly understood parameter space of dust modelling on a population scale (e.g. Bennassuti et al., 2014; Hirashita, 2015; Popping et al., 2017; Vijayan et al., 2019; Yates et al., 2024).
@@
In this work, we present a dust evolution model for incorporation into the upcoming COLIBRE suite of next-generation cosmological hydrodynamical simulations (Schaye et al. in prep., Chaikin et al. in prep.), which include cold gas physics in the multiphase ISM. We track dust via a relatively lightweight grain representation, comprising two grain sizes and three material species (one carbonaceous and two silicate subspecies). In §2 we first describe the components of our dust model (in terms of dust composition and life-cycle, with gas-dust and dust-dust interaction processes). In §3, we then briefly describe the other elements of the COLIBRE galaxy formation model, focussing on those most relevant or connected to our dust implementation. We then explore the connection between dust and ISM physics, and salient model parameters mediating this in §3.6. We present results of this model in our fiducial test simulation run in §4, focussing on dust properties and scaling relations of the low-redshift ($z=0.1$) population. We discuss our overall findings and present a summary of the paper and main conclusions in §5.

### v2

@@ Abstract
While marginal in mass terms, dust grains play an outsized role in both the physics and observation of the interstellar medium (ISM). However, explicit modelling of this ISM constituent remains uncommon in large cosmological simulations. In this work, we present a model for the life-cycle of dust in the ISM that couples to the forthcoming COLIBRE galaxy formation model, which explicitly simulates the cold ISM. We follow 6 distinct grain types: 3 chemical species, including carbon and two silicate grains, with 2 size bins each. Our dust model accounts for seeding of grains from stellar ejecta, self-consistent element-by-element metal yields and growth by accretion, grain size transfer (shattering and coagulation) and destruction of dust by thermal sputtering in the ISM. We detail the calibration of this model, particularly the use of a clumping factor, to account for unresolved gas clouds in which dust readily evolves. We present a fiducial run in a 253 cMpc3 cosmological volume that displays good agreement with observations of the cosmic evolution of dust density, as well as the $z=0$ galaxy dust mass function and dust scaling relations. We highlight known tensions between observational datasets of the dust-to-gas ratio as a function of metallicity depending on which metallicity calibrator is used; our model favours higher-normalisation metallicity calibrators, which agree with the observations within 0.1 dex for stellar masses $10^{9}\;{\rm M_{\odot}}$. We compare the grain size distribution to observations of local galaxies, and find that our simulation suggests a higher concentration of small grains, associated with more diffuse ISM and the warm-neutral medium (WNM), which both play a key role in boosting H2 content. Putting these results and modelling approaches in context, we set the stage for upcoming insights into the dusty ISM of galaxies using the COLIBRE simulations.
@@ 1 Introduction
While the interaction of stars and multi-phase gas have been a primary focus in the field of galactic astrophysics, cosmic dust, a vital piece of the puzzle of how galaxies form, has received less attention. Despite comprising a marginal fraction of the baryonic mass within galaxies, these small, solid-state particulates in the interstellar medium (ISM) play a prominent role in our understanding of the physics of galaxies and their evolution. The influence of dust belies its marginal mass contribution and is felt in a variety of ISM physics; influencing cooling and heating of gas, fostering conditions for star formation, and various dynamical effects. Dust depletes metals from the gas phase (e.g. Jenkins, 2009; De Cia et al., 2016), catalyses molecular hydrogen (${\rm H}_{2}$) formation (Cazaux & Tielens, 2002), can function as a ‘critical coolant’ in astrophysical gas (including the intra-cluster medium, the ISM of elliptical galaxies, protostellar clouds e.g. Mathews & Brighenti 2003; Montier & Giard 2004; Dopcke et al. 2011), and mediates radiative transfer through absorption and re-emission of stellar radiation (e.g. Whitaker et al., 2017; Barišić et al., 2020; Galliano et al., 2021).
@@
The dust grain size distribution, composition, and spatial configuration in galaxies is critical for understanding how radiation propagates through galaxies, but remain poorly constrained outside of the Milky Way and its satellites due to the complex interplay between dust production/destruction mechanisms and the ISM environment. The majority of dust is thought to comprise carbonaceous (C) grains and silicate-based grains (dominated by O, Si, Mg, Fe). Dust forms in stellar ejecta (e.g., supernovae, AGB stars), grows via accretion of gas-phase metals, and is destroyed by sputtering in hot plasmas, while shattering and coagulation modulate the grain size distribution in turbulent regions (e.g. Hirashita, 2012; Asano et al., 2013a). These processes drive a transition in dominant dust growth mechanisms with metallicity: at low metallicity, stellar ejecta dominate, while accretion becomes critical at higher metallicities ($Z\gtrsim 0.1\;Z_{\odot}$), explaining observed breaks in dust-to-gas vs. metallicity relations (e.g. Rémy-Ruyer et al., 2014; Hou et al., 2019).
Despite its importance, modelling dust evolution poses significant challenges due to additional computational complexity and profound uncertainties in physical processes like grain growth and destruction, often happening deep within dense, obscured media. Early modelling efforts focused on one-zone chemical evolution models (e.g. Dwek, 1998), but recent advances have allowed incorporation of dust into (cosmological) hydrodynamical simulations, including dust models using either a minimal grain size distribution (via e.g. a simple two-size approximation, Hirashita 2015, e.g.; Hou et al. 2019, e.g.; Granato et al. 2021, e.g.), or more extensive size distribution, modelling many grain sizes (e.g. McKinnon et al., 2018). The two-size approximation, explored in smoothed-particle hydrodynamics (SPH) and adaptive mesh refinement (AMR) codes (across idealised galaxy, zoom-in and even cosmological, hydrodynamical simulations), tracks separate populations of small ($r_{\rm grain}\lesssim 0.03\;{\rm\mu m}$) and large grains ($r_{\rm grain}\gtrsim 0.1\;{\rm\mu m}$) to capture size-dependent effects critical for modelling extinction curves and cooling rates, while forgoing excessive computational cost and the memory footprint associated with tracking a full size distribution per resolution element (Hirashita, 2015; Aoyama et al., 2017; Hou et al., 2019). However, spatial inhomogeneity of dust properties, evident in Milky Way extinction curve variations and observations of nearby galaxies (e.g. Gordon et al., 2003), and the need for multi-species treatments (carbon vs. silicate grains, e.g. Shivaei et al. 2020) remain important challenges (Granato et al., 2021; Dubois et al., 2024).
@@
Cosmological zoom-in simulations provide a means to self-consistently capture dust-gas co-evolution over cosmic timescales. These can follow how dust properties emerge with the enrichment of the ISM, avoiding the artificial initial conditions of isolated disc models (Gjergo et al., 2018; Dubois et al., 2024; Choban et al., 2024; Choban et al., 2025). Zoom-in simulations can afford more computationally and memory intensive representations of the ISM, allowing more sophisticated modelling of physical processes and finer spatial resolution. However, these simulations lack the statistical population sampling of larger volume simulations. Large cosmological volume simulations including dust allow galaxy dust scaling relations and distributions to be tested, providing an exacting test of dust models (e.g. Dayal et al., 2010; Bekki, 2013; McKinnon et al., 2018; Li et al., 2019). Previously, volume simulations lacked the physics to explicitly trace the cold ISM, imposing an equation of state on the ISM gas, so cannot explicitly model the cool and cold ($\ll 10^{4}~{\rm K}$) ISM that is critically important for dust evolution processes, pushing more of the relevant physics to sub-grid modelling. Given these limitations for population scales for simulations, semi-analytic models have provided insight into the potential balance of dust evolutionary processes, and have provided a means to explore the poorly understood parameter space of dust modelling on a population scale (e.g. Bennassuti et al., 2014; Hirashita, 2015; Popping et al., 2017; Vijayan et al., 2019; Yates et al., 2024).
@@
In this work, we present a dust evolution model for incorporation into the upcoming COLIBRE suite of next-generation cosmological hydrodynamical simulations (Schaye et al., 2025; Chaikin et al., 2025), which include cold gas physics in the multiphase ISM. We track dust via a relatively lightweight grain representation, comprising two grain sizes and three material species (one carbonaceous and two silicate subspecies). In §2, we first briefly describe the elements of the COLIBRE galaxy formation model outside the dust model presented here, focussing on those most relevant or connected to our dust implementation. In §3 we then describe the components of our dust model itself (in terms of dust composition and life-cycle, with gas-dust and dust-dust interaction processes). We then explore the connection between dust and ISM physics, and salient model parameters mediating this in §3.7. We present results of this model in our fiducial test simulation run in §4, focussing on dust properties and scaling relations of the low-redshift ($z=0.1$) population. We discuss our overall findings and present a summary of the paper and main conclusions in §5.
@@
2 The COLIBRE model & Simulations
[…]
2.1 Energy feedback
[…]
For the thermal part, the target heating temperature, $\Delta T_{\rm SN}(n_{\rm H})$, is an increasing function of local gas density with resolution-dependent floor and ceiling values ($6.5<\log_{10}\Delta T_{\rm SN,min}/{\rm K}<7$ and $7.5<\log_{10}\Delta T_{\rm SN,max}/{\rm K}<8$) calibrated from test runs. In the intermediate range between these extrema, $\Delta T_{\rm SN}(n_{\rm H})=10^{6.5}\;{\rm K}(n_{\rm H}/n_{\rm H,pivot})^{2/3}$. This approach mitigates overly destructive feedback and extensive bubbles in low-density gas. Finally, we adopt the statistically isotropic gas particle particle selection for energy injection of Chaikin et al. (2022), with the thermal feedback also using a variable fraction of energy per single CCSN in units of $10^{51}~\mathrm{erg}$. This is subject to a variable coupling efficiency, $f_{\rm E}$, to avoid over-cooling at high density (Crain et al., 2015). In COLIBRE this is implemented as a sigmoid in SSP birth pressure, calibrated to reproduce the $z=0$ galaxy stellar mass function and galaxy sizes (Chaikin et al., 2025).
Alongside stars, COLIBRE accounts for feedback by supermassive black holes (SMBHs). SMBH particles are first seeded for friends-of-friends halos (identified on-the-fly) in a resolution-dependent way, using an SMBH particle of fixed mass ($3\times 10^{4}\;{\rm M_{\odot}}$ at m6), for halos above a halo FoF mass ($10^{10}\;{\rm M_{\odot}}$ at m6). These particles can then accrete gas, merge and feed energy back to the surrounding media. The COLIBRE model includes two prescriptions for AGN feedback: purely thermal energy injections and a hybrid model combining thermal and kinetic jet feedback. In this work, we employ the purely thermal feedback prescription. Following similar arguments for effective feedback in CCSN, feedback is quantised to heat surrounding gas by a target temperature increment, $\Delta T_{\rm AGN}$, to mitigate numerical cooling losses. SMBHs build up an energy reservoir until enough is accrued for a single feedback event. The energy for feedback comes from gas accretion; mass energy is converted to a luminosity, that couples back to the gas with fixed efficiency ($5\%$ at m6 resolution). In COLIBRE, $\Delta T_{\rm AGN}$ is a function of black hole mass, again enforcing floor and ceiling values ($\log_{10}\Delta T_{\rm AGN,min}/{\rm K}=6.5$ and a resolution-dependent $9.5<\log_{10}\Delta T_{\rm SN,max}/{\rm K}<10$) helping to maintain a reasonable sampling and regularity of BH feedback across the mass range of AGN. Feedback is meted out to the gas particle closest to the SMBH at feedback time.
2.2 Star formation
Star formation rates are computed for eligible gas particles, which correspond to a probability of wholesale conversion into stars over a given time increment. Star formation then proceeds stochastically, sampling gas particles that are to be converted to star particles in the following timestep.
2.3 Stellar mass loss and chemical enrichment
Enrichment timesteps for star particles are designed to balance accurate sampling of stellar evolution within each SSP against computational cost. To capture the early, most active phase of enrichment, we use fine timesteps ($\leq 1$ Myr) for SSPs younger than 40 Myr. For older populations (>100>100 Myr), the time intervals after which mass transfer is performed increase proportionally to age.
2.4 COLIBRE heating and cooling without live dust
For P25 dust, it is assumed that grains are immediately destroyed at $T>10^{5}\;{\rm K}$ and do not exist in relatively diffuse gas, represented by a minimum hydrogen column density, ($N_{\rm min}=3.1\times 10^{15}\;{\rm cm}^{-2}$), such that $\mathcal{DTG}_{\rm hyb}(T>10^{5}\;{\rm K})=0$ and $\mathcal{DTG}_{\rm hyb}(N_{\rm H}\leq N_{\rm min})=0$. This corresponds to a range in volumetric hydrogen number densities at which shielding processes can occur of $n_{\rm H}>10^{-8}\;{\rm cm^{-3}}$. At high column densities ($N_{\rm H}>10^{20}\;{\rm cm}^{-2}$), a constant dust-to-metal ratio is assumed, such that [^3] As in P25, we use the $Z_{\odot}=0.0134$ value of Asplund et al. (2009) where appropriate. $\mathcal{DTG}_{\rm hyb}\propto Z_{\rm gas}/{\rm Z_{\odot}}$. For intermediate column densities ($3.1\times 10^{15}<N_{\rm H}/{\rm cm}^{-2}<10^{20}$), an additional taper is introduced to smooth the transition between the metallicity-dependent $\mathcal{DTG}_{\rm hyb}(N_{\rm H}=10^{20}\;{\rm cm}^{-2})\propto Z_{\rm gas}/{\rm Z_{\odot}}$ and $\mathcal{DTG}_{\rm hyb}(N_{\rm H}=3.1\times 10^{15})=0$, assuming a power law in column density with index $\alpha=1.4$. This is chosen based on the Kennicutt-Schmidt relation (Kennicutt, 1998), and assuming a scaling between the star formation rate density and dust production. At high density and solar metallicity, the $\mathcal{DTG}$ is normalised to a representative MW value of $\mathcal{DTG}_{\rm MW}=6.6\times 10^{-3}$.
where $\nu_{\rm HI}$ and $n_{\rm HI}$ are the thermal velocity and the physical number density of atomic hydrogen [^4] We note that dust properties assumed by the cooling module do not make use of the density boost introduced in §3.8, $\epsilon_{\rm H_{2}}$ is the recombination efficiency of molecular hydrogen, $S_{\rm H}$ is a dimensionless sticking coefficient and $n_{\rm d}\sigma_{\rm d}$ is the number density of dust grains times the grain cross section, i.e. the total cross section per unit volume. While $S_{\rm H}$ and $\sigma_{\rm d}$ are temperature-dependent, a constant dust temperature of 10 K is assumed [^5] The H2 formation rate was shown to be approximately constant for $6\;{\rm K}\lessapprox T_{\rm dust}\lessapprox 50\;{\rm K}$ in fig. A1 of Richings et al. (2014a).. It is then the $n_{\rm d}\sigma_{\rm d}$ term that depends on the local mass fraction and size distribution of grains, with the other terms depending on the material properties and local gas properties. Dust grains can also contribute to catalysis of ion recombination via available dust surfaces (see e.g. Weingartner & Draine, 2001), and transfer energy between the dust and gas (Richings et al., 2014b).
The influence of dust can also be felt indirectly by processing radiative energy via extinction. A particularly important example of this is photoelectric heating by dust grains; absorption of FUV photons by dust produces energetic free electrons that contribute to a heating of the gas component. In CHIMES, the volumetric heating rate, $\Gamma_{\rm PE}$, follows that of Wolfire et al. (2003) (their equations 19 and 20666This is an update of what is described in Richings et al. (2014a), described by Richings & Faucher-Giguère (2018).), and is given by
With the grain cross-section per unit volume the relevant quantity for all of these processes, we can modulate them using the masses and sizes of grains traced through our live dust model. We describe their coupling in §3.7. A final dust-sensitive aspect to consider is the role of depletion. The same chemical elements that constitute dust also play an important cooling role in the ISM when existing in the gas-phase, particularly through metal-line cooling. These processes are handled via per-ion cooling rates tabulated for COLIBRE and based upon Oppenheimer & Schaye (2013) (P25 equation 19888Updated from Ploeckinger & Schaye (2020)). These rates are based on gas-phase metals, so it is necessary to assume a per-element depletion alongside the $\mathcal{DTG}_{\rm hyb}$, in lieu of a live dust model. These assume depletion in the fixed relative proportions of the Milky Way as inferred by Jenkins (2009). We discuss how we then adjust these rates for the case of live dust in §3.7, consistent with the depletions we calculate for our dust model (§3.9).
2.5 Simulation Runs
Table 1: Simulation variations used in this work. All runs represent a 253 Mpc3 cosmological box, initialised with 3763 gas particles of mass $m_{\rm g}=1.8\times 10^{6}\;{\rm M_{\odot}}$ (and 4 times as many dark matter particles, with mass $m_{\rm DM}=2.4\times 10^{6}\;{\rm M_{\odot}}$). We adopt ‘m6’ as shorthand for this resolution (from the COLIBRE parlance). Simulations are run using the COLIBRE galaxy formation model (Schaye et al. in prep.), using the swift code (Schaller et al., 2024, see section 2.5). The columns from left to right indicate: 1) the simulation name, 2) if the dust model is coupled to the cooling (see § 2.4), 3) if the model includes small grains rather than large, if we include processes of 4) accretion, 5) destruction 6) or size transfer, 7) the maximum $C$ value 8) if $C$ has a fixed value 9) the size representing the small grain bin 10) the chemical group assumed for silicates and $1$) the diffusion constant (§ 3.6).
1) 2) 3) 4) 5) 6) 7) 8) 9) 10) 11)
(a) [µm]
Fiducial 100. 0.01 Olivine 0.01
FidUncoupled 100. 0.01 Olivine 0.01
SeedOnly n/a 0.01 Olivine 0.01
NoDestruction 100. 0.01 Olivine 0.01
OneSize 100. 0.01 Olivine 0.01
MaxC10 10. 0.01 Olivine 0.01
constC30 30. 0.01 Olivine 0.01
NoC 1. 0.01 Olivine 0.01
PyroSil 100. 0.01 Pyroxene 0.01
SmallerGrains 100. 0.005 Olivine 0.01
NoDiff 100. 0.01 Olivine 0
LoDiff 100. 0.01 Olivine 0.001
HiDiff 100. 0.01 Olivine 0.1
(a) i.e. the clumping factor
Prior to describing our dust model in detail (§3, next), we describe the simulations used in this work and our choice of parameter variations. Throughout this work we assume the $3\times 2$ pt cosmology, combining all data and constraints presented with the DES Y3 results (Abbott et al., 2022); $h=0.681$, $\Omega_{\rm m}=0.306$, $\Omega_{\rm b}=0.0486$, $\sigma_{8}=0.807$, $n_{\rm s}=0.967$.
[…]
We identify collapsed structures using the friends-of-friends (FoF) algorithm post processed using the publically available HBT-HERONS subhalo finder described by Forouhar Moreno et al. (2025) to assign particles to self-bound substructures. This builds on the HBT+ code for $N$-body simulations (Han et al., 2018), with critical improvements for the hydrodynamical context, considering the histories and hierarchical formation of subhalos. In particular, the HBT-HERONS code exhibits more consistent allocation of mass between satellite and central galaxies during close passes, as compared to various other phase-space halo finders (Chandro-Gómez et al., 2025). The subhalo locations and, optionally, the particles are then taken up by the SOAP tool (McGibbon et al., 2025), computing a wide range of aggregated properties using a range of 3D and projected apertures of varying physical size.
The simulations run for analysis in this work and their defining parameters are listed in Table 1. The Fiducial run uses the full dust model described in §3, and also includes the effects of “coupling” our dust model to gas cooling, heating and phase transition processes, detailed in §3.7. Our runs incorporate dust coupling by default (excepting SeedOnly where dust fractions are too unphysical), with an uncoupled counterpart to isolate the influence of the dust evolution model on the cooling, e.g. FidUncoupled. The SeedOnly and NoDestruction simulations isolate evolutionary processes. Size transfer effects are tested by LargeGrainsOnly. A number of runs are dedicated to exploring the influence of parameters; particularly a subgrid clumping factor, $C$, (described in §3.8; NoC, ConstC30 and MaxC10) and the turbulent diffusion constant, $C_{\rm d}$ (see §3.6; NoDiff, LoDiff and HiDiff). We also present runs exploring pyroxene silicate chemistry, PyroSil, and the representative size of small grains, SmallerGrains. The use of coupled cooling physics by default means we see the impact of these dust model changes on the physics and resultant galaxy populations. We will explore the most pertinent variations alongside our primary low-redshift results in §4, as well as presenting some additional variations in Appendix A.